OpenStudy (anonymous):

Two violin strings are tuned to the same frequency, 277 Hz. The tension in one string is then decreased by 1.92 percent. What will be the beat frequency heard when the two strings are played together?

OpenStudy (anonymous):

please help the assignment is due in 30mins :(

OpenStudy (anonymous):

ok, what expression are you using for frequency as a function of tension?

OpenStudy (anonymous):

f=sqrt T?

OpenStudy (anonymous):

pretty much, all else being constant and for the purpose of the problem so:

OpenStudy (anonymous):

\[\frac{ f _{1} }{f _{2}} = \frac{ \sqrt{.9808T} }{ \sqrt{T}}\]

OpenStudy (anonymous):

Is that all that youre given?

OpenStudy (anonymous):

\[{ f _{1} } = \sqrt{.9808} {f _{2}}\]

OpenStudy (anonymous):

my f2 is 277?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so find the freq. of the loosened string.

OpenStudy (anonymous):

the beat frequency is the difference between the two.

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

thanks alot sweet heart

OpenStudy (anonymous):

no problem, sugar.